# Factor Risk Budgeting and Beyond

### v1

@@ Abstract
Portfolio optimization methods have evolved significantly since Markowitz introduced the mean-variance framework in 1952. While the theoretical appeal of this approach is undeniable, its practical implementation poses important challenges, primarily revolving around the intricate task of estimating expected returns. As a result, practitioners and scholars have explored alternative methods that prioritize risk management and diversification. One such approach is Risk Budgeting, where portfolio risk is allocated among assets according to predefined risk budgets. The effectiveness of Risk Budgeting in achieving true diversification can, however, be questioned, given that asset returns are often influenced by a small number of risk factors. From this perspective, one question arises: is it possible to allocate risk at the factor level using the Risk Budgeting approach? This paper introduces a comprehensive framework to address this question by introducing risk measures directly associated with risk factor exposures and demonstrating the desirable mathematical properties of these risk measures, making them suitable for optimization. We also propose a framework to find the portfolio that effectively balances the risk contributions from both assets and factors. Leveraging standard stochastic algorithms, our framework enables the use of a wide range of risk measures.
Keywords: factor models, factor investing, portfolio optimization, risk budgeting, risk measures, stochastic algorithms
@@ 1 Introduction
Portfolio selection is a two-stage procedure, as emphasized in Markowitz’ seminal 1952 paper [20]. The first stage involves forming beliefs about the future evolution of asset returns, and the second stage concerns portfolio construction based on these beliefs. Modern portfolio theory provides fundamental results for the second stage, by mathematically formalizing the benefits of the ubiquitous concept of diversification and proposing a clear investment objective for investors, based on mean-variance analysis. However, the complexity of the first stage often leads to erroneous beliefs on expected returns that Markowitz’ portfolios excessively reflect in asset weightings (see [3] and [24]).
@@
Two main lines of research have therefore coexisted to improve the portfolio selection process in academia and in the asset management industry.
@@
On the one hand, a considerable amount of effort has been devoted to improving the second stage, by reducing the sensitivity to input parameters and obtaining practically more plausible portfolios. Some methods have attempted to retain expected returns within the picture, such as the Black-Litterman model [4] and entropy pooling [22], while other methods omit them fully (or partially) and only involve the covariance matrix in the portfolio selection process – so-called risk-based methods. Among risk-based methods, there are several fundamental approaches based on different rationales. [^1] Interestingly, most of them boil down mathematically to minimizing the portfolio risk under a specific constraint with different parameters (see [26]). Minimum Variance (MV) portfolio is one of them, corresponding to the portfolio with the minimum risk (as measured by variance) among all the possible fully invested portfolios. Another approach is that of the Most-Diversified portfolio, as defined in [9], which seeks to maximize the ratio between the weighted sum of the volatilities of individual assets and the total volatility of the portfolio. Equal Risk Contribution (or Risk Parity) is another risk-based approach that differs from the previous two in motivation. Rather than focusing on optimizing a final statistic of the portfolio, it aims to allocate portfolio risk equally across its assets. Risk Budgeting (RB) is a generalization of this approach, which allows the user to choose the target risk contribution of each asset by setting asset-specific risk budgets. Risk Budgeting has become a popular method, particularly over the last two decades, thanks to some interesting theoretical properties (see [5], [6] and [28]).
@@
On the other hand, with regard to the first stage of portfolio selection, the quest for a better understanding of the true drivers of asset returns and the risk-return relationship dates back to the 1960s, when the Capital Asset Pricing Model (CAPM) was independently developed by Sharpe, Lintner, Mossin and Treynor (see [19, 25, 31, 34]). The CAPM establishes a linear relationship between expected returns of risky assets and that of the market portfolio, assuming that asset returns are governed by a single factor. The arbitrage pricing theory (APT) developed in [30] extends this original approach to a multi-factor framework, without identifying the risk factors. The Fama-French three-factor model proposed in [12] is one of the most famous factor models, expanding the CAPM by introducing size and value factors on top of the market factor. There are other prominent models that later introduced new factors like momentum, profitability and investments (see [7] and [13]). Over the years, the number of factors proposed by academics and practitioners has grown considerably, giving rise to the phenomenon of the factor zoo, as stated by [10]. However, a recent study argues that most of these factors are redundant (see [32]).
@@
In the context of a factor model, exploring and controlling exposure to underlying factors become crucial. For example, [16] explores portfolios designed to achieve specific factor exposures, highlighting the existence of infinitely many portfolios meeting the desired factor exposures. In [11], the effectiveness of the equally-weighted factor portfolio is compared to alternative factor-allocation schemes. Although risk-based approaches have been explored, in particular the feasibility of Risk Budgeting to obtain diversified factor-focused portfolios, it should be noted that, with the exception of a few papers dealing with factor risk contributions in the context of correlated factors (see [21] and [29]), most papers employ factor risk budgeting portfolios on uncorrelated factors. This is often achieved by using techniques like principal / independent component analysis or Minimum-Torsion Bets to decorrelate existing factors (see [1], [17] and [23]).
This paper aims to precisely define the problem of Factor Risk Budgeting by providing the necessary tools to calculate portfolios that take into account factor risk contributions for a wide range of risk measures.
@@
In Section 2, we outline the classical Risk Budgeting problem and present an efficient method to find Risk Budgeting portfolios. In Section 3, we present linear factor models and the concept of factor exposures. Then, we introduce a factor risk measure which maps factor exposures to risk. We show that these factor risk measures possess all the necessary mathematical properties to be used for Risk Budgeting. Section 4 precisely defines the Factor Risk Budgeting problem. We propose a framework in the same section that seeks a balance between classical Risk Budgeting and Factor Risk Budgeting, considering both asset and factor risk contributions. Section 5 presents two case studies where we construct equity portfolios under the BARRA risk model and multi-asset portfolios in the context of a macroeconomic factor model to enhance our understanding of Factor Risk Budgeting portfolios and its variants.
@@ 2.1 Notations and main concepts
In asset management, the risk is associated with a portfolio made of multiple assets. Therefore, we consider a financial universe of $d$ assets whose joint returns are given by a random vector $X=(X_{1},\ldots,X_{d})^{\prime}$ where $(1,X_{1},\ldots,X_{d})$ is assumed to be linearly independent. Given a vector of asset exposures denoted by $y\in\mathbb{R}^{d}$, the loss associated with $y$ is the random variable $-y^{\prime}X$ and the risk is $\rho(-y^{\prime}X)$. [^3] The vector $y$ can be interpreted as a portfolio where short positions and leverage are allowed. This leads to a function $\mathcal{R}_{\rho,X}$ that we call an asset exposure risk measure associated with $\rho$ (and $X$):
@@
$\mathcal{R}_{\rho,X}$ is hereafter denoted by $\mathcal{R}$ when there is no ambiguity. We assume that $\mathcal{R}$ is continuous on $\mathbb{R}^{d}$ and continuously differentiable on $\mathbb{R}^{d}\setminus\{\mathbf{0}\}$. We also assume that $\forall y\in\mathbb{R}^{d}\setminus\{\mathbf{0}\}$, $\mathcal{R}(y)>0$, i.e. the risk of a non-zero financial position is always positive. We finally make the assumption that $\mathcal{R}$ is additive solely under scaled asset exposures.
@@
$\mathcal{R}(\theta)=\sum_{i=1}^{d}\theta_{i}\partial_{i}\mathcal{R}(\theta)$
where $\partial_{i}\mathcal{R}(\theta)$ is the partial derivative of $\mathcal{R}$ with respect to the variable $\theta_{i}$ at point $\theta$. In the above sum the $i^{\text{th}}$ term is called the risk contribution of asset $i$ to the total risk of the portfolio. The concept of risk contribution stands as a fundamental pillar within the Risk Budgeting framework.
@@ 2.2 Formulation of the Risk Budgeting problem and associated theoretical results
[…]
@@
[…]
Questions that naturally arise from the above definition concern the existence and uniqueness of a solution. The theorem below proved in [8] states the existence of a unique solution to the Risk Budgeting problem.
@@
[…]
The importance of Theorem 1 goes beyond confirming the existence of a unique solution. Rather than directly engaging with the system of non-linear equations that characterize the original problem given by Definition 1, this theorem highlights the possibility of approaching the Risk Budgeting problem through unconstrained convex minimization problems. This latter approach has clear benefits in terms of mathematical properties and computational efficiency. Furthermore, the choice of the function $g$ is free, making the approach versatile as we illustrate in the next section.
@@ 2.3 Computation of Risk Budgeting portfolios
Early research on Risk Budgeting and most applications have focused on the case where the co-movements of asset returns are measured by their covariance matrix $\Sigma$ and the risk of a portfolio with asset exposures $y$ is measured by its volatility, i.e. $\mathcal{R}(y):=\sqrt{y^{\prime}\Sigma y}$ ($\mathcal{R}$ satisfies our assumptions whenever $\Sigma$ is invertible – an assumption that we consider throughout the paper). For a given $\Sigma$, the volatility of the portfolio and its gradient can easily be computed. Consequently, the Risk Budgeting problem for volatility can be solved efficiently using a gradient descent procedure, by choosing $g(x)=x^{2}$ and minimizing over (+*$\mathbb{(R_{+}^{*})}^{d}$ the function
@@
However, the scope of risk measures used in portfolio management goes well beyond volatility, considering the inherent asymmetry and heavy-tailed nature of asset return distributions. Accordingly, the literature on Risk Budgeting has sometimes touched upon more general risk measures. For instance, in [18], Risk Budgeting portfolios are constructed for Expected Shortfall. [^4] The Expected Shortfall at level $\alpha\in(0,1)$ of a continuous random variable (loss) $Z\in L^{1}$ with positive density is defined by $\mathbb{E}[Z|Z\geq\text{VaR}_{\alpha}(Z)]$ where $\mathbb{P}(Z\geq\text{VaR}_{\alpha}(Z))=1-\alpha$. In order to employ the aforementioned gradient descent procedure effectively, they make a specific assumption about asset returns. Precisely, they assume that asset returns follow a mixture of two multivariate Gaussian distributions – a setting in which Expected Shortfall exhibits a semi-analytic formula (see [8] for an extension to Student mixtures). In cases where there are no parametric assumptions on asset returns or when dealing with more complex risk measures, gradient computation often necessitates approximation through computationally intensive methods (see [14]).
@@
In fact, the set of risk measures that enable the above stochastic path for computation of Risk Budgeting portfolios is quite broad. All deviation measures and spectral risk measures (as formalized in [8]) fall within this framework. Volatility and Expected Shortfall are the two most popular members of these classes, respectively. For volatility, i.e. $\rho(Z)=\min_{\zeta\in\mathbb{R}}\sqrt{{\mathbb{E}}\left[(Z-\zeta)^{2}\right]}$, the $\text{RB}_{b}$ problem becomes, [^5] Of course, using a stochastic gradient descent algorithm in the case of volatility is not the best option. A gradient descent on Eq. (1) is more efficient. upon choosing $g:x\mapsto x^{2}$,
@@
It is essential to highlight that this approach is suited to a broad spectrum of asset return distributions. Consequently, to some extent, this approach can be considered universal with respect to asset return distributions. Furthermore, it exhibits a high level of versatility, when it comes to using different risk measures. For many risk measures, a stochastic gradient descent algorithm based on the above ideas is a great option to effectively tackle Risk Budgeting problems.

### v2

# Asset and Factor Risk Budgeting: A Balanced Approach

@@ Abstract
Portfolio optimization methods have evolved significantly since Markowitz introduced the mean-variance framework in 1952. While the theoretical appeal of this approach is undeniable, its practical implementation poses important challenges, primarily revolving around the intricate task of estimating expected returns. As a result, practitioners and scholars have explored alternative methods that prioritize risk management and diversification. One such approach is Risk Budgeting, where portfolio risk is allocated among assets according to predefined risk budgets. The effectiveness of Risk Budgeting in achieving true diversification can, however, be questioned, given that asset returns are often influenced by a small number of risk factors. From this perspective, one question arises: is it possible to allocate risk at the factor level using the Risk Budgeting approach? First, we introduce a comprehensive framework to address this question by introducing risk measures directly associated with risk factor exposures and demonstrating the desirable mathematical properties of these risk measures, making them suitable for optimization. Then, we propose a novel framework to find portfolios that effectively balance the risk contributions from both assets and factors. Leveraging standard stochastic algorithms, our framework enables the use of a wide range of risk measures to construct diversified portfolios.
Keywords: diversification, factor models, factor investing, portfolio optimization, risk budgeting, risk measures, stochastic algorithms
@@ 1 Introduction
In Markowitz’ seminal 1952 paper [24], portfolio selection is described as a two-stage procedure. The first stage involves forming beliefs about the future evolution of asset returns, and the second stage concerns portfolio construction based on these beliefs. Modern portfolio theory provides fundamental results for the second stage, by mathematically formalizing the benefits of the ubiquitous concept of diversification and proposing a clear investment objective for investors, based on mean-variance analysis. However, the complexity of the first stage often leads to erroneous beliefs on expected returns that Markowitz’ portfolios excessively reflect in asset weightings (see [4] and [28]).
@@
Two main lines of research have therefore coexisted to improve the portfolio selection process in the academic literature and in the asset management industry.
@@
On the one hand, a considerable amount of effort has been devoted to improving the second stage, by reducing the sensitivity to input parameters and obtaining practically more plausible portfolios. Some methods have attempted to retain expected returns within the picture, such as the Black-Litterman model [5] and entropy pooling [26], while other methods omit them fully (or partially) and only involve the covariance matrix in the portfolio selection process – so-called risk-based methods. Among risk-based methods, there are several fundamental approaches based on different rationales. [^1] Interestingly, most of them boil down mathematically to minimizing the portfolio risk under a specific constraint with different parameters (see [17] and [30]). Minimum Variance (MV) portfolio is one of them, corresponding to the portfolio with the minimum risk (as measured by variance) among all the possible fully invested portfolios. Another approach is that of the Most-Diversified portfolio, as defined in [10], which seeks to maximize the ratio between the weighted sum of the volatilities of individual assets and the total volatility of the portfolio. Equal Risk Contribution (or Risk Parity) is another risk-based approach that differs from the previous two in motivation. Rather than focusing on optimizing a final statistic of the portfolio, it aims to allocate portfolio risk equally across its assets. Risk Budgeting (RB) is a generalization of this approach, which allows the user to choose the target risk contribution of each asset by setting asset-specific risk budgets. Risk Budgeting has become a popular method, particularly over the last two decades, thanks to some interesting theoretical properties (see [6], [7] and [32]).
@@
On the other hand, with regard to the first stage of portfolio selection, the quest for a better understanding of the true drivers of asset returns and the risk-return relationship dates back to the 1960s, when the Capital Asset Pricing Model (CAPM) was independently developed by Sharpe, Lintner, Mossin and Treynor (see [23, 29, 35, 38]). The CAPM establishes a linear relationship between the expected excess returns of risky assets and the expected excess return of the market portfolio, resulting in a single-factor model for asset returns. The arbitrage pricing theory (APT) developed in [34] extends this original approach to a multi-factor framework, without identifying the risk factors. The Fama-French three-factor model proposed in [14] is one of the most famous factor models, expanding the CAPM by introducing size and value factors on top of the market factor. There are other prominent models that later introduced new factors like momentum, profitability, quality, investments, etc. (see for instance [8] and [15]). Over the years, the number of factors proposed by academics and practitioners has grown considerably, giving rise to the phenomenon of the factor zoo, as stated by [11]. However, a recent study argues that most of these factors are redundant (see [36]).
@@
In the context of a factor model, exploring and controlling exposure to underlying factors become crucial. For example, [19] explores portfolios designed to achieve specific factor exposures, highlighting the existence of infinitely many portfolios meeting the desired factor exposures. In [13], the effectiveness of the equally-weighted factor portfolio is compared to alternative factor-allocation schemes. Although risk-based approaches have been explored, in particular the feasibility of Risk Budgeting to obtain diversified factor-focused portfolios, it should be noted that, with the exception of a few papers dealing with factor risk contributions in the context of correlated factors (see [25] and [33]), most papers consider factor risk budgeting portfolios on uncorrelated factors. This is often achieved by using techniques like principal / independent component analysis or Minimum-Torsion Bets to decorrelate existing factors (see [1], [21] and [27]).
The primary objective of this paper is to propose a rigorous mathematical framework for constructing diversified portfolios that effectively mitigate the concentration of risk at both the asset and factor levels, especially in the presence of dependent or correlated risk factors. In a setting where asset returns are driven by specific factors, the central challenge for portfolio managers is to manage exposures to these factors effectively. Our first contribution addresses this challenge by introducing a specific formalization through a factor risk measure – a mapping from factor exposures to risk – that enables the selection of a unique vector of asset exposures. We demonstrate the financial viability and essential mathematical properties of the proposed factor risk measure, which are crucial for its effective use in risk management and portfolio optimization. These findings and ideas might be of interest beyond risk budgeting techniques for factor investing in general.
Additionally, we show that our factor risk measure leads to a formal definition and solution of the Factor Risk Budgeting problem, which involves allocating portfolio risk across chosen risk factors according to predefined risk budgets for each factor. This represents the second contribution of our work. Furthermore, we present an optimization framework that integrates the principles of Factor Risk Budgeting with classical Risk Budgeting to design portfolios that account for both asset and factor risk contributions. To the best of our knowledge, this integrated framework represents a novel contribution to the asset management literature and aims to serve as a practical methodology for achieving diversification. Finally, we highlight the stochastic formulation of the proposed optimization problem, which allows portfolio managers to efficiently work with a wide range of risk measures when quantifying portfolio risk.
@@
In Section 2, we outline the classical Risk Budgeting problem and present an efficient method to find Risk Budgeting portfolios. In Section 3, we present linear factor models and the concept of factor exposures. Then, we introduce factor risk measures which map factor exposures to risk. We show that these factor risk measures possess all the necessary mathematical properties to be used for risk management and portfolio optimization, especially for Risk Budgeting. Section 4 precisely defines the Factor Risk Budgeting problem. In Section 5, we propose the Asset-Factor Risk Budgeting framework that seeks a balance between classical Risk Budgeting and Factor Risk Budgeting, considering both asset and factor risk contributions. Section 6 presents two case studies where we construct equity portfolios under the BARRA risk model and cross-asset portfolios in the context of a macroeconomic factor model to exemplify the use of Factor Risk Budgeting and Asset-Factor Risk Budgeting portfolios.
@@ 2.1 Notations and main concepts
In asset management, the risk is associated with a portfolio made of multiple assets. Therefore, we consider a financial universe of $d$ assets whose joint returns are given by a random vector $X=(X_{1},\ldots,X_{d})^{\prime}$ where $(1,X_{1},\ldots,X_{d})$ is assumed to be linearly independent. Given a vector of asset exposures denoted by $y\in\mathbb{R}^{d}$, the loss associated with $y$ is the random variable $-y^{\prime}X$ and the risk is $\rho(-y^{\prime}X)$. [^3] The vector $y$ can be interpreted as a portfolio where short positions and leverage are allowed. This leads to a function $\mathcal{R}_{\rho,X}$ that we call an asset exposure risk measure associated with $\rho$ and $X$:
@@
$\mathcal{R}_{\rho,X}$ is hereafter denoted by $\mathcal{R}$ when there is no ambiguity. We assume that $\mathcal{R}$ is continuous on $\mathbb{R}^{d}$ and continuously differentiable on $\mathbb{R}^{d}\setminus\{\mathbf{0}\}$. [^4] The risk measure $\rho$ and the distribution of $X$ are assumed to be chosen to ensure that the mentioned properties for $\mathcal{R}$ are satisfied. We also assume that $\forall y\in\mathbb{R}^{d}\setminus\{\mathbf{0}\}$, $\mathcal{R}(y)>0$, i.e. the risk of a non-zero financial position is always positive. We finally make the assumption that $\mathcal{R}$ is additive solely under scaled asset exposures:
@@
$\mathcal{R}(\theta)=\sum_{i=1}^{d}\theta_{i}\partial_{i}\mathcal{R}(\theta),$
where $\partial_{i}\mathcal{R}(\theta)$ is the partial derivative of $\mathcal{R}$ with respect to the variable $\theta_{i}$ at point $\theta$. In the above sum, the $i^{\text{th}}$ term is called the risk contribution of asset $i$ to the total risk of the portfolio. The concept of risk contribution is a fundamental notion in the Risk Budgeting framework.
@@ 2.2 Formulation of the Risk Budgeting problem and associated theoretical results
Let $b\in\Delta^{>0}_{d}$ be a vector of risk budgets. We say that a vector of weights $\theta\in\Delta^{>0}_{d}$ solves the Risk Budgeting problem $\text{RB}_{b}$ if and only if
@@
This is the classical definition of Risk Budgeting where only strictly long-only portfolios are taken into account. If we remove this constraint, there exist multiple solutions in $\Delta_{d}$ to $\text{RB}_{b}$ (see [3] for a detailed analysis in the case of volatility).
Questions that naturally arise from the above definition concern the existence and uniqueness of a solution. The constructive theorem that follows, proved in [9], states the existence of a unique solution to the Risk Budgeting problem.
@@
There exists a unique solution $\theta\in\Delta^{>0}_{d}$ of $\text{RB}_{b}$ and $\theta=\frac{y^{*}}{\sum_{i=1}^{d}y^{*}_{i}}$.
The importance of Theorem 1 goes beyond confirming the existence of a unique solution. Rather than directly engaging with the system of non-linear equations that characterize the original problem given by Definition 1, this theorem highlights the possibility of approaching the Risk Budgeting problem through unconstrained convex minimization problems. The latter technique has clear benefits in terms of mathematical properties and computational efficiency. Furthermore, the choice of the function $g$ is free, making the approach versatile, as we illustrate in the next section.
@@ 2.3 Computation of Risk Budgeting portfolios
Early research on Risk Budgeting and most applications have focused on the case where the co-movements of asset returns are measured by their covariance matrix $\Sigma$ and the risk of a portfolio with asset exposures $y$ is measured by its volatility, i.e. $\mathcal{R}(y):=\sqrt{y^{\prime}\Sigma y}$ ($\mathcal{R}$ satisfies our assumptions whenever $\Sigma$ is invertible – an assumption that we consider throughout the paper). For a given $\Sigma$, the volatility of the portfolio and its gradient can easily be computed. Consequently, the Risk Budgeting problem for volatility can be solved efficiently using a gradient descent procedure, by choosing $g(x)=x^{2}$ and minimizing over $\mathbb{(R_{+}^{*})}^{d}$ the function
@@
However, the scope of risk measures used in portfolio management goes well beyond volatility and includes a wide variety of symmetric and asymmetric risk measures. In [2], the authors consider Risk Parity when the risk measure is chosen to be mean absolute deviation around the mean and provide a comprehensive mathematical analysis of the specific problem. Considering the inherent asymmetry and heavy-tailed nature of asset return distributions, the literature on Risk Budgeting has also touched upon other asymmetric risk measures. For instance, in [22], Risk Budgeting portfolios are constructed for Expected Shortfall. [^5] The Expected Shortfall at level $\alpha\in(0,1)$ of a continuous random variable (loss) $Z\in L^{1}$ with positive density is defined by $\mathbb{E}[Z|Z\geq\text{VaR}_{\alpha}(Z)]$ where $\mathbb{P}(Z\geq\text{VaR}_{\alpha}(Z))=1-\alpha$. In order to employ the aforementioned gradient descent procedure effectively, they make a specific assumption about asset returns. Precisely, they assume that asset returns follow a mixture of two multivariate Gaussian distributions – a setting in which Expected Shortfall exhibits a semi-analytic formula (see [9] for an extension to Student mixtures). In cases where there are no parametric assumptions on asset returns or when dealing with more complex risk measures, gradient computation often necessitates approximations through computationally intensive methods (see [16]). [^6] In [20], the authors tackle the Risk Budgeting problem for dynamic risk measures.
@@
In fact, the set of risk measures that enable the above stochastic path for the computation of Risk Budgeting portfolios is quite broad. All deviation measures and spectral risk measures (as formalized in [9]) fall within this framework. Volatility and Expected Shortfall are the two most popular members of these classes, respectively. For volatility, i.e. $\rho(Z)=\min_{\zeta\in\mathbb{R}}\sqrt{{\mathbb{E}}\left[(Z-\zeta)^{2}\right]}$, the $\text{RB}_{b}$ problem becomes, [^7] Of course, using a stochastic gradient descent algorithm in the case of volatility is not the best option. A gradient descent on Eq. (1) is more efficient. upon choosing $g:x\mapsto x^{2}$,
@@
It is essential to highlight that this approach is suited to a broad spectrum of asset return distributions. Furthermore, it exhibits a high level of versatility, when it comes to using different risk measures. For many risk measures, a stochastic gradient descent algorithm based on the above ideas is a great option to effectively tackle Risk Budgeting problems. [^8] In [12], the authors propose using cutting planes algorithm for the specific case of Expected Shortfall. Beyond volatility and Expected Shortfall, some other examples of risk measures covered by the stochastic framework are mean absolute deviation around the median, power spectral risk measure and variantile [39].